OpenStudy (anonymous):

An athlete can run 100m in 11 seconds. Calculate the athlete's speed in- metre's per second, and kilomstres per hour

OpenStudy (anonymous):

m/s = 100/11 km/hr=100/11*1/1000*3600

OpenStudy (anonymous):

100m/11s=9.0909091m/s in km/hr: (9.0909091m/s)/1000=0.00909091km; (0.00909091)*3600=32.72724km/hr; (1hr=3600s)